OpenStudy (tay1lynn):

Marcie solved the following inequality, and her work is shown below: −2(x − 5) − 12 ≤ 4 + 6(x + 3) −2x + 10 − 12 ≤ 4 + 6x + 18 −2x − 2 ≤ 6x + 22 −8x ≤ 24 x ≤ −3 What mistake did Marcie make in solving the inequality? She subtracted 6x from both sides when she should have added 2x. She added 2 to both sides when she should have subtracted 22. When dividing by −8, she did not change the ≤ to ≥. She did not make a mistake.

OpenStudy (jocelm):

Doesn't look like she made a mistake to me.

OpenStudy (maddyhall15):

This is correct.

OpenStudy (tay1lynn):

ok thank you so much

OpenStudy (zzr0ck3r):

When you divide or multiply by a negative you need to flip the inequality sign.

OpenStudy (zzr0ck3r):

@JoCelM and @maddyhall15 and @tay1lynn

OpenStudy (zzr0ck3r):

\(−8x ≤ 24\implies x\ge \dfrac{24}{-8}\implies x\ge -3\)

OpenStudy (zzr0ck3r):

Understand?

OpenStudy (tay1lynn):

Yes I do now thank you

OpenStudy (jocelm):

Wow I didn't catch that! Good job.